Chapter 1 Proton Structure Chapter 2 Spin Physics Chapter 3 Production of Hadrons, Jets and Photons Chapter 4 Heavy-Flavour Production Chapter 5 Elastic and Diffractive ep Scattering Chapter 6 New Physics at HERA Chapter 7 Plans for HERA-II and Beyond